2 5 7 11 14 19 23 29 ?

Answer by Edwin McCravy(20056)

Sometimes if you'll look at every other term, starting with the first
term you'll see one pattern, and starting with the second term, you'll
see a different pattern.    

2 5 7 11 14 19 23 29 ?

Look at the sequence of the odd numbered terms

2   7    14    23    ?

 2 + 5 =  7
 7 + 7 = 14
14 + 9 = 23  See the pattern? So the next term will be:
23 +11 = 34

Look at the sequence of the even numbered terms

  5   11    19    29 ?

 5 + 6 = 11
11 + 8 = 19
19 +10 = 29    See the pattern? So the next even numbered term will be:
29 +12 = 41 

So the sequence goes:

2 5 7 11 14 19 23 29 34 41

Edwin